OpenStudy (anonymous):

i need to show work on this one, find the value of x in 5x+6=3x-6

OpenStudy (anonymous):

so we need to rewrite this so our x's are on one side and our constants are on the other 5x-3x=-6-6 2x=-12 x=-6